OpenStudy (anonymous):

Which set of line segments could create a right triangle? 15, 30, 35 15, 36, 39 15, 20, 29 5, 15, 30

OpenStudy (anonymous):

Hello, This question requires the Pythagorean theorem, ever heard of it? It looks like this: \[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

We apply the theorem by taking the smallest sides and assigning them to the variables \(a\) and \(b\). When variables \(a\) and \(b\) are equal to \(c\) then the set of line segments are possible to create a right triangle. @Bballplayer48 do you follow so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So A @Algorithmic

OpenStudy (anonymous):

No \(A\) would not work. \[15^2 + 30^2 \ne 35^2\]

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

D would?

OpenStudy (anonymous):

No... Are you guessing? It is really simple to solve this problem provided you follow my instructions above...

OpenStudy (anonymous):

Oh I understand

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Bingo!

OpenStudy (anonymous):

You add both a and b when you get the exponet then you find the squareroot

OpenStudy (anonymous):

15 to the second power + 36 to the second power = 1521 and then the square root is 39

OpenStudy (anonymous):

B's work if anyone else is interested: \[15^3 + 36^2 = 1521\] \[39^2 = 1521\] \[1521 = 1521\] Hence, B is the correct answer. Peace.